# Supermassive black holes swallow stellar objects at high rates: from Little Red Dots to Black Hole Stars

### v2

@@ VIII Formation of the black-hole star
Accretion-disk sizes inferred from reverberation mapping are typically parsec-scale. The Eddington feedback region is set by requiring a unit Thomson optical depth,
@@
using $\rho_{g}=\sigma^{6}G^{-3}M_{g}^{-2}$ and stellar mass form factors (Diener et al., 1995). This defines a dense central core at radii $\sim 1$$10\,{\rm AU}$, consistent with CLOUDY models of inverted LRD spectra and inferred gas densities $n_{H}\sim 10^{8}\,{\rm cm^{-3}}$ (Taylor et al., 2025).
Observations and theory favor Comptonization-regulated feedback (Gilli et al., 2022): momentum-driven regulation dominates at high redshift, while energy-driven feedback with momentum boosting becomes important at low redshift, enhancing early star formation and later driving massive outflows (Silk et al., 2024; Costa et al., 2014). Specifically,
  • Momentum-driven feedback: $M_{\rm BH}\propto\sigma^{4}(\sigma_{T}/m_{p})/(\pi G^{2})$; inverse-Compton cooling yields insufficient momentum by a factor $\sim 10$, but can trigger positive star-formation feedback (Silk and Nusser, 2010).
  • Energy-driven feedback: $M_{\rm SMBH}=\alpha\sigma^{5}(\sigma_{T}/m_{p})/(\pi\eta cG^{2})$, providing the required momentum boosting and reproducing the observed $M_{\rm BH}$$\sigma$ relation.
@@ IX Conclusions
Compact-object captures produce thousands of detectable LISA sources within $z<6$ and generate a stochastic GW background, dominated by WD captures at deciHz frequencies.
Stellar feeding is therefore an inevitable byproduct of dense, rapidly growing high-redshift nuclei; insufficient for SMBH growth, but observable through extreme TDEs and GWs, providing a direct probe of black-hole–star systems and early SMBH assembly.
@@
We thank Mesut Çalışkan, Marco Chiaberge, Muryel Guolo, Francesco Iacovelli, Mitchell Karmen, Julian Krolik, Colin Norman, and David Pereñiguez for discussions and comments. $K.$ is supported by NSF Grants No. AST-2307146, PHY-2513337, PHY-090003, and PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International, by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167, and Onassis Foundation Scholarship (ID: F ZT 041-1/2023-2024).
Appendix A Extended tables of loss-cone rates
This Appendix presents capture loss-cone rates and associated uncertainty for each stellar type at the five different epochs onto a $10^{5}\,M_{\odot}$ (Table 2), $10^{6}\,M_{\odot}$ (Table 3), and $10^{7}\,M_{\odot}$ (Table 4) SMBH.
$t\ (\rm Myr)$ 1 10 100 1,000 10,000
@@ Appendix A Extended tables of loss-cone rates
Table 2: Capture loss-cone rates onto $M_{\rm SMBH}=10^{5}\,M_{\odot}$ in $\rm yr^{-1}$ at evolution time $t$ for different stellar types. Rates of naked cores are zero. Error bars are standard deviations estimated over ten realizations. Total number of stellar objects within the influence radius is $N_{\star}=301,550$. Abbreviations: MS (main sequence); WD (white dwarf); NS (neutron star); BH (black hole).
$t\ (\rm Myr)$ 1 10 100 1,000 10,000

### v3

@@ VIII Formation of the black hole star
Accretion-disk sizes inferred from reverberation mapping are typically parsec scale. The Eddington feedback region is set by requiring a unit Thomson optical depth,
@@
using $\rho_{g}=\sigma^{6}G^{-3}M_{g}^{-2}$ and stellar-mass form factors (Diener et al., 1995). This defines a dense central core at radii $\sim 1$$10\,{\rm au}$, consistent with CLOUDY models of inverted LRD spectra and inferred gas densities $n_{\rm H}\sim 10^{8}\,{\rm cm^{-3}}$ (Taylor et al., 2025).
Observations and theory favor Comptonization-regulated feedback (Gilli et al., 2022): momentum-driven regulation dominates at high redshift, while energy-driven feedback with momentum boosting becomes important at low redshift, enhancing early star formation and later driving massive outflows (Silk et al., 2024; Costa et al., 2014).
@@ IX Conclusions
Compact-object captures produce thousands of detectable LISA sources within $z<6$ and generate a stochastic GW background, dominated by WD captures at dHz frequencies.
Stellar feeding is therefore an inevitable byproduct of dense, rapidly growing high-redshift nuclei; insufficient for SMBH growth, but observable through extreme TDEs and GWs, providing a direct probe of BH–star systems and early SMBH assembly.
@@
We thank Mesut Çalışkan, Marco Chiaberge, Muryel Guolo, Francesco Iacovelli, Mitchell Karmen, Julian Krolik, Colin Norman, David Pereñiguez, and the anonymous referee for discussions and comments. $K.$ is supported by NSF grant Nos. AST-2307146, PHY-2513337, PHY-090003, and PHY-20043, by NASA grant No. 21-ATP21-0010, by John Templeton Foundation grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International, by Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167, and by the Onassis Foundation Scholarship (ID: F ZT 041-1/2023-2024).
Extended tables of loss-cone rates
This Appendix presents capture loss-cone rates and associated uncertainties for each stellar type at the five different epochs onto a $10^{5}$ (Table 2), $10^{6}$ (Table 3), and $10^{7}\,M_{\odot}$ (Table 4) SMBH.
stellar class $t\ (\rm Myr)$
1 10 100 1000 10,000
@@
Table 2: Capture loss-cone rates onto $M_{\rm SMBH}=10^{5}\,M_{\odot}$ in $\rm yr^{-1}$ at evolution time $t$ for different stellar types. Rates of naked cores are zero. Error bars are standard deviations estimated over 10 realizations. Total number of stellar objects within the influence radius is $N_{\star}=301,550$. Abbreviations: MS (main sequence); WD (white dwarf); NS (neutron star); BH (black hole).
stellar class $t\ (\rm Myr)$
1 10 100 1000 10,000